find the measure of an angle whose complement is 25 less than half of its supplement.
Answer by mananth(12270)

let angle be x
xomplement = 90-x
supplement = 180-x
...
90-x = 1/2 * (180-x) -25
..
90-x = (180-x -50)/2
2(90-x)=130-x
180-2x=130-x
180-130=2x-x
50=x the angle
50 deg.